OpenStudy (anonymous):

find the exact value: sin (105 degrees) cos (15 degrees). Somebody explain me the steps please? Medal to helper.

OpenStudy (anonymous):

look first we should split the (105) into a familiar angles (90 , 45 , 30 ,60 ,etc...) so since 105 = 45 + 60 then sin(105)=sin(45+60) and the same with the cosine. then solve it as any sum of 2 angles ;)

OpenStudy (anonymous):

but look out for the angle (15) you should use "the equation of the half of the angle" : \[\cos(\frac{ x }{ 2 }) = \sqrt{\frac{ 1+cosx }{ 2 }}\]